Consider a sample of H2 and N2 at 1.00 bar and 25oC with mole fraction (X) of N2=0.25 Calculate the number of molecules m-3 for H2 and N2
Answer #1

The sample consists of H2 and N2

Pressure = 1 bar

Temperature = 25 C = 298 K

Volume = 1 m^3 = 1000 L

We will use the ideal gas equation to calculate the moles of mixture of gas

PV = nRT

n = number of moles = Pressure X volume / R X temperature

R = gas constant = 0.08314 bar L / mole K

Moles = 1 X 1000 / 0.08314 X 298 = 40.36 moles

The mole fraction of N2 = 0.25

So moles of N2 = Mole fraction x total moles = 0.25 x 40.36 = 10.09 mole

Number of molecules of N2 = 10.09 X 6.023 X 10^23 = 60.77 X 10^23

Moles of H2 = Mole fraction x total moles = 0.75 X 40.36 = 30.27

Number of molecules of H2 = 30.27 X 6.023 X 10^23 = 182 .31 X 10^23